find the equilibrium level income and rate of interest from the following information

C=40+0.8Yd, 110-5r, T=50, G=50, Md=0.2y-5r, Ms=100


1
Expert's answer
2021-07-21T14:20:58-0400

"Y = C + I + G = 40 + 0.8(Y - 50) + 110 - 5r + 50 = 0.8Y + 160 - 5r,"

0.2Y = 160 - 5r.............(I)

Ms = Md, so:

0.2Y - 5r = 100,

0.2Y = 100 + 5r,...….....(ii)

160 - 5r = 100 + 5r,

10r = 60,

r = 6%

Hence from equation (ii)

0.2Y = 100+5r

0.2Y = 100+(5×6)


"Y = \frac{100 + (5×6)}{0.2 }= 650."